rewrite the equation in standard form with integer coefficients.
2y=-2x+6
y=3/4x+1/2

Answer by drk(1908)

Standard form requires two things: (1) all variable on the same side and (2) no fractions.
--
2y=-2x+6
Here we need variables on the same side so, we get
2x + 2y = 6.
Since there are no fractions, we are done.
--
y=3/4x+1/2
Here we have fractions. Multiply by 4, the common denominator. We get
4y = 3x + 2
Now put all variables on the same side. We get
-3x + 4y = 2.
Some teachers are picky about this and require x to be positive.